# The Surprising Soupability of Documents in State Space Models

### v1

@@ Abstract
We investigate whether hidden states from Structured State Space Models (SSMs) can be merged post-hoc to support downstream reasoning. Inspired by model souping, we propose a strategy where documents are encoded independently and their representations are pooled—via simple operations like averaging—into a single context state. This approach, which we call document souping, enables modular encoding and reuse without reprocessing the full input for each query. We finetune Mamba2 models to produce soupable representations and find that they support multi-hop QA, sparse retrieval, and long-document reasoning with strong accuracy. On HotpotQA, souping ten independently encoded documents nearly matches the performance of a cross-encoder trained on the same inputs.
@@ 1 Introduction
Many real-world NLP tasks—such as multi-document question answering, scientific summarization, and legal analysis—require reasoning over entire corpora, not just individual long documents. These tasks demand flexible integration of information spread across sources, as well as the ability to update, prune, or recombine subsets of the input. Yet today’s language models remain poorly suited for this kind of modular document reasoning. Transformer-based models [21] suffer from $\mathcal{O}(L^{2})$ attention costs and memory scaling, making full-corpus encoding expensive and inflexible.
Structured State Space Models (SSMs) offer a compelling alternative. Recent architectures such as Mamba [8] and Mamba2 [6] process long sequences in linear time, compressing them into fixed-size hidden states that support efficient downstream decoding. This makes them well-suited for long-context settings. However, most current pipelines treat corpora as monolithic inputs—concatenating documents into a single sequence before encoding. This approach is brittle and inefficient: modifying even a single document requires re-encoding the entire corpus, and joint encoding eliminates opportunities for reuse across tasks or queries.
Figure 1: Computation Graphs for Corpus Encoding. Top: In traditional concatenation-based encoding, all documents $\{d_{1},\dots,d_{k}\}$, the query $q$, and answer $a$ are flattened into a single input sequence and processed end-to-end by an SSM. This requires joint re-encoding for every change to the input. Bottom: In our proposed state souping approach, each document $d_{i}$ is encoded independently by a shared SSM, producing per-document hidden states $\{h_{1},\dots,h_{k}\}$ which are pooled into a single representation $h_{\text{soup}}$ (e.g., via sum or average). This pooled state is then used, alongside the query $q$, to drive downstream prediction. The design supports parallel encoding, modular reuse, and post-hoc corpus composition.
In this work, we explore whether it is possible to modularize document representation in SSMs through a strategy inspired by model souping—a technique that averages parameters across finetuned checkpoints to combine capabilities [25]. We ask: if SSMs encode individual documents into fixed-length hidden states, can those representations be merged post-hoc—for example, by averaging—while still supporting downstream tasks like question answering? This capability would be especially valuable in retrieval-augmented generation, where cross-encoder approaches must re-encode the top-$k$ retrieved documents for every new query. In contrast, if document states are soupable, one could pre-encode the full corpus once, cache each document’s hidden state, and dynamically pool a subset at inference time—enabling scalable, query-specific reconfiguration without re-encoding.
We operationalize this idea as corpus encoding via state souping. Each document is independently encoded by a shared SSM to produce layerwise hidden states. These are then pooled using simple commutative operators (e.g., average, sum, max) into a single fixed-size representation, which conditions the decoder alongside the query. We define soupability as the property of document encodings that allows them to be pooled in this way while preserving the information needed for downstream reasoning.
Surprisingly, we find that Mamba2 encoders produce representations with strong soupability: even without joint encoding, a finetuned decoder can learn to interpret pooled document states effectively. In this setting, the encoder remains frozen, and only the decoder adapts—indicating that the individual document encodings already retain task-relevant information that can be meaningfully composed through pooling. With full encoder-decoder finetuning, performance improves further, as the encoder learns to produce even more mergeable representations.
This architecture enables a new approach to multi-document reasoning. For example, on HotpotQA [28]—a benchmark requiring multi-hop reasoning across documents—souping ten independently encoded documents achieves nearly the same accuracy as a standard cross-encoder, where all documents are jointly re-encoded with the query. More broadly, we find that corpus encoding via state souping supports single- and multi-hop QA, sparse retrieval, and long-document understanding—often matching or surpassing traditional joint encoding approaches.
Crucially, this design unlocks a new inference workflow: corpora can be encoded once, cached as hidden states, and later reassembled dynamically for downstream tasks. This supports post-hoc, retrieval-based reasoning at scale, without the need for repeated full-sequence processing.
@@
We present a systematic investigation into the soupability of document representations in Structured State Space Models (SSMs) using Mamba2-2.7B [6] and Mamba2-8B [22]. Our contributions are: (1) We introduce document souping, a mechanism for merging per-document hidden states using lightweight pooling operators and injecting the result into an SSM decoder. (2) We demonstrate that, after training, Mamba2 models with soup-based representations can support multi-hop reasoning on HotpotQA, achieving performance comparable to or exceeding joint encoding. (3) We analyze soupability across diverse tasks and conditions—including long-document QA (RACE[14]/QuALITY[16]), sparse retrieval (RULER[11]/NIAH), and document count generalization—showing that souped states preserve both local and relational information. (4) We compare pooling strategies and show that soupability is robust to operator choice, with no-norm averaging working well across settings.
@@ 2 Methods
We investigate when document representations from Structured State Space Models (SSMs) can be merged post-hoc—using simple commutative operations like averaging or summation—while still preserving the information needed for downstream tasks. We refer to this property as soupability.
@@ 2.1 Background and Notation
Structured State Space Models (SSMs) are a class of sequence models that compress long inputs into fixed-size hidden representations through linear recurrence mechanisms. Unlike attention-based models, SSMs offer subquadratic compute and memory scaling, making them well suited for long-context settings.
We denote the encoder component of an SSM as a function $\mathrm{SSM}_{\theta}$, which maps a document $d$ to a sequence of layer-wise hidden states $\{h^{(1)},\dots,h^{(L)}\}$. Each $h^{(l)}\in\mathbb{R}^{d}$ summarizes the input up to layer $l$, and $L$ denotes the total number of layers. In standard pipelines, this encoder is applied to the entire input sequence jointly, and the resulting states are passed to a decoder for generation or prediction.
In this work, we leverage the layerwise structure of SSMs to explore whether documents can be encoded independently and later recombined in a modular fashion. Our central question is whether these hidden states—computed separately per document—can be pooled into a single representation suitable for downstream reasoning. We refer to this approach as corpus encoding via state souping.
@@ 2.2 Corpus Encoding via State Souping
Our proposed corpus encoding strategy is illustrated in Figure 1 and formalized in Appendix A.3. Given a set of documents $\{d_{1},\dots,d_{k}\}$ and a query $q$, each document is passed independently through a shared SSM encoder, yielding a set of hidden states $\{h_{1}^{(l)},\dots,h_{k}^{(l)}\}$ at each layer $l$. These are then combined using a commutative pooling operator—typically elementwise average, sum, or max—resulting in a single pooled state $h_{\text{soup}}^{(l)}$ for each layer.
@@ 2.3 Evaluation Dimensions
Model capacity concerns the architectural and training properties that affect soupability. We ask: Are pretrained SSMs inherently soupable, or must they be finetuned to interpret pooled states? Does soupability improve with model size or hidden state dimensionality? And how well do models generalize across soup sizes—for example, when asked to merge more documents at test time than during training?
@@
Method Test on $2$ gold + $(n{-}2)$ distractors
@@
Table 1: HotpotQA performance (Exact Match / F1) for Mamba2-8B models trained on 5 documents (2 gold + 3 distractors) and evaluated on $n$ documents, each with 2 gold and $(n{-}2)$ distractors. We compare pretrained models, decoder-only finetuning, and encoder-decoder finetuning across different pooling strategies (average, sum, max), with optional normalization applied before and/or after aggregation. Underlined entries indicate evaluations where the number of test-time documents matches the training configuration. We observe that decoder-only finetuning improves performance by learning to interpret fixed souped states, and full encoder-decoder finetuning yields the best results by also learning to produce mergeable representations. Across all configurations, simple averaging without normalization emerges as the most stable and effective aggregation method.
@@ 3 Experiments
We study multi-document question answering in two distinct regimes. For single-hop QA, we use the QA subset of the RULER dataset [11], which augments SQuAD-style [19] questions for long-context evaluation. Each question is answerable from a single gold document placed within a large set of distractors, allowing us to isolate how well the model can identify and preserve localized information in souped representations. For multi-hop QA, we turn to HotpotQA [28], a benchmark requiring compositional reasoning across multiple Wikipedia paragraphs. Each question demands integration of evidence from at least two documents, providing a direct test of whether souped hidden states preserve the relational structure needed for multi-hop reasoning.
We also evaluate on long-document QA tasks, where inputs are single extended narratives rather than disjoint documents. In this setting, we train on the RACE dataset [14], which consists of relatively short educational passages, and test on both RACE and the validation portion of QuALITY [16], which features much longer and more complex narratives. This setup allows us to assess whether document segmentation and aggregation via souping generalizes from short-form to long-form content. Our findings support earlier observations from the QuALITY paper that models trained on RACE can transfer effectively.
Finally, we evaluate sparse retrieval using the multikey-2 subset of the RULER dataset. In this setting, each document line contains a unique identifier paired with a corresponding value, and the model is tasked with memorizing all such mappings. At inference time, it receives a query specifying one of the identifiers and must output the correct value. The full input includes many such mappings, requiring the model to store a dense set of key-value pairs and retrieve the correct one after aggregation. This task challenges the model’s ability to preserve fine-grained, instance-specific information across multiple independently encoded documents.
Together, these tasks span diverse input formats and reasoning types, enabling a broad and controlled evaluation of state souping. Unless otherwise noted, all experiments are conducted using Mamba2-8B. Results for the 2.7B model are reported in Appendix B. In general, we observe that the larger model consistently achieves stronger performance, particularly in settings requiring multi-hop reasoning or generalization across segment length.
Table 2: HotpotQA results for Mamba2-8B evaluated on 2 gold documents plus $(n{-}2)$ distractors. We compare concat-based and soup-based training across a range of finetuning configurations: QA-only (no documents), 2-gold only, and 2-gold + distractors. Each model is evaluated at multiple soup sizes, and underlined entries mark evaluations that match the training number of documents.
Method Test on $2$ gold + $(n{-}2)$ distractors
@@
[…]
Full Finetuned on $n=0\ gold+0\ distractors$ (QA-only)
@@
Full Finetuned on $n=2\ gold+0\ distractors$ documents
@@
[…]
Full Finetuned on $n=2\ gold+3\ distractors$ documents
@@
[…]
Full Finetuned on $n=2\ gold+5\ distractors$ documents
@@
[…]
Full Finetuned on $n=2\ gold+8\ distractors$ documents
@@ 3.1 Experimental Setup
3.1.1 Training Configurations
To ensure reproducibility, we fix all random seeds to 42 and enable deterministic=True. Optimization follows the setup from [8], using AdamW with $\beta_{1}=0.9$, $\beta_{2}=0.95$, gradient clipping at 1.0, and a cosine decay schedule with 10% linear warm-up. The 2.7B model is finetuned with a learning rate of $5\times 10^{-5}$, and the 8B model with $1\times 10^{-5}$. All experiments are trained for a single epoch, with one exception: the single-hop QA setting. For single-hop QA, we constructed a dataset of approximately 6,000 examples by augmenting SQuAD-style questions using the RULER framework. Of these, 5,000 examples were used for training. Due to the smaller dataset size, one epoch of training was insufficient for convergence, so we trained for 3 epochs in this setting.
Gradient accumulation is set to 120 steps111128 for the 2.7B HotpotQA configuration.. For distributed training, we use DeepSpeed ZeRO-2 [18], which allows each GPU to hold the full model and facilitates the extraction of complete hidden states for the souping mechanism. Experiments were mainly conducted on clusters with H100 and H200 GPUs; Full training and implementation details are provided in Appendix A.
3.1.2 Evaluation Metrics
We evaluate model performance using standard metrics for both extractive and multiple-choice question answering. For extractive QA tasks such as HotpotQA and RULER, we report exact match (EM), which measures the percentage of predictions that exactly match any ground-truth answer span, and F1 score, which reflects the token-level overlap between the predicted and reference answers. The F1 score is defined as the harmonic mean of precision and recall:

### v2

# Studying the Soupability of Documents in State Space Models

@@ Abstract
We investigate whether hidden states from Structured State Space Models (SSMs) can be merged post hoc to support downstream reasoning. Inspired by model souping, we study document souping, a strategy where documents are encoded independently, and their representations are pooled, via simple operations like averaging, into a single context state. This approach enables modular encoding and reuse without reprocessing the full input for each query. We demonstrate that finetuned Mamba2 models with souped representations achieve competitive or superior performance across multi-hop QA, sparse retrieval, and long-document reasoning tasks compared to the standard monolithic encoding approach. For example, on the RACE and QuALITY benchmarks for long document question answering, this method substantially outperforms a traditional concatenation approach. Crucially, this modular design scales to hundreds of documents (we test up to 256), while delivering substantial savings in inference cost, unlocking new possibilities for large-scale corpus reasoning.
Machine Learning, ICML
Figure 1: Computation Graphs for Corpus Encoding. Top: In traditional concatenation-based encoding, all documents $\{d_{1},\dots,d_{k}\}$, the query $q$, and answer $a$ are flattened into a single input sequence and processed end-to-end by an SSM. This requires joint re-encoding for every change to the input. Bottom: In the state-souping approach we study, each document $d_{i}$ is encoded independently by a shared SSM, producing per document hidden states $\{h_{1},\dots,h_{k}\}$ which are pooled into a single representation $h_{\text{soup}}$ (e.g., via sum or average). This pooled state is then used, alongside the query $q$, to drive downstream prediction. The design supports parallel encoding, modular reuse, and post hoc corpus composition.
@@ 1 Introduction
Many real-world NLP tasks, such as multi-document question answering, scientific summarization, and legal analysis, require reasoning over entire corpora rather than individual long documents. These tasks demand flexible integration of information distributed across sources, as well as the ability to dynamically update, prune, or recombine input subsets. Yet today’s language models remain poorly suited for this kind of modular document reasoning.
Transformer-based models (Vaswani et al., 2017), despite their success, face prohibitive $\mathcal{O}(L^{2})$ attention costs that make full corpus encoding expensive and inflexible. Structured State Space Models (SSMs) offer a promising alternative: architectures like Mamba (Gu and Dao, 2024) and Mamba2 (Dao and Gu, 2024) process sequences in linear time, compressing them into fixed-length hidden states. The linear recurrence central to these models provides the core intuition for our work: we hypothesize that linear operations on these states, such as averaging, will produce meaningful composite representations. This architectural feature is key, as we later show empirically that it makes SSMs uniquely suited for the state-souping setup we study, an approach that proves ineffective in standard Transformers. However, even these efficient architectures are typically deployed using a monolithic encoding strategy: concatenating all documents into a single sequence before processing. This approach inherits a critical weakness: any modification to the corpus, even changing a single document, requires re-encoding the entire input from scratch and prevents the reuse of document representations. This re-encoding requirement becomes prohibitive at scale, processing hundreds or thousands of documents repeatedly for each query.
This brittleness suggests a fundamental mismatch: we want modular, reusable representations, yet mainstream approaches still rely on monolithic, use-once encodings. In this work, we explore an alternative inspired by model souping (Wortsman et al., 2022), which merges finetuned checkpoints by averaging their parameters. Recent work has begun to push this idea to hidden states: pióro2024statesoupincontextskill show that in-context skills can be stored as recurrent states, retrieved, and linearly mixed, while Liu et al. (2025) derives a theoretically well-motivated operator for composing SSM states that approximates a permutation-invariant version of full concatenation and includes an initial investigation into souped-state representations. However, these approaches leave open a key question: can factual knowledge spread across long documents and corpora be reliably souped and composed to support real multi-document reasoning, and what failure modes arise when one attempts to do so?
We address this question from a complementary, empirically driven perspective. If SSMs encode individual documents into fixed-length hidden states, can those representations be merged post hoc while still supporting downstream tasks such as multi-hop QA and long-document reading comprehension? We refer to this property as soupability: the ability of independently encoded document representations to be pooled while preserving the information needed for multi-document reasoning. Our aim is to test souping capabilities for realistic multi-document and long-context settings, analyze when it fails, and what training and pooling regimes make it effective in practice.
To study soupability concretely, we operationalize it as corpus encoding via state souping, illustrated in Figure 1. Each document is independently encoded by a shared SSM to produce per-document hidden states, which are then pooled using simple commutative operators (e.g., averaging) into a single representation that conditions the decoder alongside the query. This construction exposes both the strengths and weaknesses of souping: naive pooling generally does not work out of the box, but with appropriate finetuning and pooling choices we find that factual content in larger documents and corpora can be reliably souped and composed, yielding performance competitive with joint encoding on challenging multi-hop benchmarks. At the same time, the approach retains a key advantage of modularity: documents can be encoded once, cached, and later combined in arbitrary subsets, enabling efficient corpus updates and dynamic retrieval without re-encoding.
We use Mamba2 as our representative SSM because its core ingredients are shared across modern SSM families, including S4 (Gu et al., 2022) and S5 (Smith et al., 2023) variants. Moreover, it is the only large-scale SSM without a hybrid architecture. This approach depends on the linear recurrent scan that produces encoder hidden states, not on details unique to Mamba2. The properties that enable souping, namely independent encoding, linear aggregation, and conditioning on a pooled state, should therefore carry over to other SSM architectures with comparable state representations.
We find that Mamba2 encodes representations with strong soupability. With full encoder-decoder finetuning, our souping-based architecture supports multi-hop QA, sparse retrieval, and long document understanding, often matching or surpassing traditional joint encoding approaches. This design unlocks a new inference workflow where corpora can be encoded once, cached, and reassembled dynamically, enabling scalable, retrieval-based reasoning without repeated full sequence processing.
@@
We systematically investigate the soupability of document representations in SSMs using Mamba2-2.7B (Dao and Gu, 2024) and Mamba2-8B (Waleffe et al., 2024). Our contributions are: (1) We provide the first systematic empirical study of document souping, its behavior, and limitations. (2) We demonstrate that finetuned Mamba2 models with souped representations can support multi-hop reasoning on HotpotQA (Yang et al., 2018), with performance comparable to joint encoding. (3) We analyze soupability across diverse tasks, including long document QA (RACE (Lai et al., 2017) and QuALITY (Pang et al., 2022)) and sparse retrieval. (4) We compare pooling strategies and show that soupability is robust to operator choice. (5) We show that this modular souping approach is uniquely suited to SSMs, as an analogous technique fails in standard Transformer architectures.
@@ 2 Methods
We investigate when document representations from Structured State Space Models (SSMs) can be merged post hoc, using simple commutative operations like averaging or summation, while still preserving the information needed for downstream tasks. We refer to this property as soupability.
@@ 2.1 Background and Notation
Structured State-Space Models (SSMs) are a class of sequence models that compress long inputs into fixed-size hidden representations through linear recurrence mechanisms. Unlike attention-based models, SSMs offer subquadratic compute and memory scaling, making them well-suited for long-context settings.
We denote the encoder component of an SSM as a function $\mathrm{SSM}_{\theta}$, which maps a document $d$ to a sequence of layer-wise hidden states $\{h^{(1)},\dots,h^{(L)}\}$. Each $h^{(l)}\in\mathbb{R}^{d}$ summarizes the input up to layer $l$, and $L$ denotes the total number of layers. This encoder is applied to the entire input sequence jointly, and the resulting states are passed to a decoder for generation or prediction.
In this work, we leverage the layerwise structure of SSMs to explore whether documents can be encoded independently and later recombined in a modular fashion. Our central question is whether these hidden states, computed separately per document, can be pooled into a single representation suitable for downstream reasoning. Given their foundation in linear recurrence, we hypothesize that a linear combination of these states could form a meaningful composite representation. We refer to this approach as corpus encoding via state souping.
@@ 2.2 Corpus Encoding via State Souping
The corpus encoding strategy we study is illustrated in Figure 1 and formalized in Appendix A.3. Given a set of documents $\{d_{1},\dots,d_{k}\}$ and a query $q$, each document is passed independently through a shared SSM encoder, yielding a set of hidden states $\{h_{1}^{(l)},\dots,h_{k}^{(l)}\}$ at each layer $l$. These are then combined using a commutative pooling operator, typically elementwise average, sum, or max, resulting in a single pooled state $h_{\text{soup}}^{(l)}$ for each layer.
@@ 2.3 Evaluation Dimensions
Model capacity concerns the architectural and training properties that affect soupability. We ask: Are pretrained SSMs inherently soupable, or must they be finetuned to interpret pooled states? Does soupability improve with model size or hidden state dimensionality? And how well do models generalize across soup sizes, for example, when asked to merge more documents at test time than during training?
@@
Table 1: HotpotQA performance (Exact Match / F1) for Mamba2-8B models trained on 5 documents (2 gold + 3 distractors) and evaluated on $n$ documents, each with 2 gold and (n2)(n{-}2) distractors. We compare pretrained models, decoder-only finetuning, and encoder-decoder finetuning across different pooling strategies (average, sum, max), with optional normalization applied before and/or after aggregation. Underlined entries indicate evaluations where the number of test-time documents matches the training configuration.
Method Test on $2$ gold + (n2)(n{-}2) distractors
@@ 3 Experiments
We conduct a comprehensive set of experiments to evaluate document souping across a range of tasks and settings. This section outlines our experimental design, detailing the datasets, baselines, and implementation specifics used in our analysis.
3.1 Tasks and Datasets
@@
Multi-Doc QA
We study multi-document question answering in two distinct regimes. For single-hop QA, we use the QA subset of the RULER dataset (Hsieh et al., 2024), which augments SQuAD-style (Rajpurkar et al., 2016) questions for long-context evaluation. Each question is answerable from a single gold document placed within a large set of distractors, allowing us to isolate how well the model can identify and preserve localized information in souped representations. For multi-hop QA, we turn to HotpotQA (Yang et al., 2018), a benchmark requiring compositional reasoning across multiple Wikipedia paragraphs. Each question demands the integration of evidence from at least two documents among a set of distractors, providing a direct test of whether souped hidden states preserve the relational structure needed for multi-hop reasoning. This setting mirrors a retrieval-augmented workflow with a small number of relevant documents among distractors, allowing us to evaluate integration quality as the retrieved set size increases.
Long Doc QA
We also evaluate on long-document QA tasks, where inputs are single extended narratives rather than disjoint documents. In this setting, we train on the RACE dataset (Lai et al., 2017), which consists of relatively short educational passages, and test on both RACE and the validation portion of QuALITY (Pang et al., 2022), which features much longer and more complex narratives. This setup allows us to assess whether document segmentation and aggregation via souping generalizes from short-form to long-form content. Our findings support earlier observations from the QuALITY paper that models trained on RACE can transfer effectively.
Sparse Retrieval
Finally, we evaluate sparse retrieval using the NIAH multikey-2 subset of the RULER dataset. In this setting, each document line contains a unique identifier paired with a corresponding value, and the model is tasked with memorizing all such mappings. At inference time, it receives a query specifying one of the identifiers and must output the correct value. The full input includes many such mappings, requiring the model to store a dense set of key-value pairs and retrieve the correct one after aggregation. This task challenges the model’s ability to preserve fine-grained, instance-specific information across multiple independently encoded documents.
Table 2: HotpotQA results for Mamba2-8B evaluated on 2 gold documents plus (n2)(n{-}2) distractors. We compare concat-based and soup-based training across a range of finetuning configurations: QA-only (no documents), 2-gold only, and 2-gold + distractors. Each model is evaluated at multiple soup sizes, and underlined entries mark evaluations that match the training number of documents.
Method Test on $2$ gold + (n2)(n{-}2) distractors
@@
Soup w/ Average 8.7 / 12.7 5.3 / 8.3 3.7 / 6.4 2.6 / 5.0 2.2 / 4.5 2.3 / 4.5 2.0 / 4.0 1.7 / 3.9 1.7 / 3.6
Full Finetuned on n=0gold+0distractorsn=0\ gold+0\ distractors (QA-only)
@@
Full Finetuned on n=2gold+0distractorsn=2\ gold+0\ distractors documents
@@
Soup w/ Average 57.1 / 70.9 51.1 / 64.4 46.7 / 59.7 43.0 / 56.0 39.2 / 52.1 36.2 / 48.5 34.2 / 46.0 32.2 / 44.0 30.6 / 42.2
Full Finetuned on n=2gold+3distractorsn=2\ gold+3\ distractors documents
@@
Full Finetuned on n=2gold+5distractorsn=2\ gold+5\ distractors documents
@@
Soup w/ Average 55.5 / 69.4 52.2 / 66.0 50.3 / 63.9 48.1 / 61.5 46.2 / 59.5 45.0 / 57.6 43.2 / 55.8 41.8 / 54.2 40.3 / 52.8
Full Finetuned on n=2gold+8distractorsn=2\ gold+8\ distractors documents
@@ 3.2 Experimental Setup
Training Configurations
All experiments use Mamba2-8B, as we observe that it consistently achieves stronger performance than the 2.7B model (see Appendix. B for comparison), particularly in settings requiring multi-hop reasoning or generalization across segment length. For most experiments, we train for a single epoch. The only exception is the single-hop QA task, which, due to its smaller dataset size, is trained for three epochs to ensure convergence. All other hyperparameters, along with specifics on the compute resources are detailed in Appendix A.4.
Evaluation Metrics
We evaluate model performance using standard metrics for both extractive and multiple-choice question answering. For extractive QA tasks such as HotpotQA and RULER, we report exact match (EM), which measures the percentage of predictions that exactly match any ground-truth answer span, and F1 score, which reflects the token-level overlap between the predicted and reference answers. For multiple-choice QA tasks, including RACE and QuALITY, we use multiple-choice accuracy as the evaluation metric. In this setting, the model outputs a distribution over the four answer options (A, B, C, D), and we compute the log-probability of each choice. The predicted answer corresponds to the option with the highest score, and accuracy is measured as the proportion of correctly selected answers.
3.3 Baselines and Reference Setting
We compare souping against several SSM baselines that represent different levels of supervision and context integration. First, we include a pretrained baseline: an off-the-shelf Mamba2 model used without any task-specific finetuning, which serves as a lower bound since the model has not been adapted to the QA tasks. Next, we consider a QA-only finetuning baseline trained only on $(q,a)$ pairs without access to any document context during training. Although this configuration does not support long-context integration, it shows how much performance is attributable to the context. Finally, we evaluate a concat-based finetuning baseline, where the model is trained end-to-end on full input sequences, i.e., $(d_{1},\dots,d_{k},q,a)$, concatenated into a single flat input. This setting provides strong supervision by exposing the model to all relevant documents in a joint context window and serves as our primary reference point when assessing the performance of souping-based alternatives. Our aim is to demonstrate that, with appropriate finetuning, souping-based models can outperform context-free baselines and approach the performance of the concat-based reference model.
Justifying the Focus on SSMs vs. Transformers
A natural question is whether document souping can be extended to Transformer architectures for the sake of comparison. However, a direct application is conceptually challenging, as Transformers do not compress inputs into a fixed-size hidden state in the same manner as SSMs. The closest analogue to an SSM’s state is the key-value (KV) cache. Crucially though, the KV cache’s size is proportional to the input sequence length, making the aggregation of caches from documents of varying lengths non-trivial. To investigate empirically, we designed an experiment to test a cache souping approach. To address the size mismatch, all documents were padded to a uniform length before being encoded independently. We implemented a cache souping variant for LLaMA 3 8B on HotpotQA, averaging the KV caches from ten documents (see Appendix A.5.1 for full experimental details). The experiment confirmed this approach is ineffective for Transformers: while standard finetuning on concatenated documents achieved 48.50 EM and 62.81 F1, cache souping yielded only 0.53 EM and 8.18 F1 given the same amount of data. This significant performance degradation supports our hypothesis that the fixed-size, recurrent state of SSMs is a critical architectural feature for this style of modular recombination. Consequently, we focus our main experiments on SSM baselines.
Relation to RAG